NTJ Distributors Ltd has has an annual demand for an airport metal detector of 1400 units. The cost of a typical detector is $400. Carrying cost is estimated to be 20% of the unit cost and the ordering cost is $25 per order. If Tricia Joseph, the owner, orders in quantities of 300, he can get a 4% discount on the cost of the detectors. Should the company accept the discount?


1
Expert's answer
2022-03-20T06:43:27-0400

If the owner orders in quantities multiple of 300 they should order 1500 units, which will be cost $400*1500 = $600,000. In that case, the discount constitutes $600,000*0.04 = $24,000.

The carrying cost for 100 detectors will be $400*100*0.2 = $8,000 < $24,000. So the company should accept the discount.